### v1

@@ 5.3 Results
Evolving hyper-parameters. We monitor the evolution of the interpolation coefficients during training of GPT-2 (125M) model on OpenWebText for various choices of initialization of $\beta_{1},\beta_{2}$ (Figures  6, 6, and 6).
@@
First, we observe that sub-optimal initialization of hyper-parameters (as in Figure 6) results in a more significant update of the coefficients, even though the hyper-learning rates are the same (except for parameters $\beta_{1},\beta_{2}$). This suggests that optimizer state with $\beta_{1,0}=0.9,\beta_{2,0}=0.95$ is itself close to a local minimum. A closer inspection of individual coefficients yields further insight into how MADA tries to accelerate training. First, in Figures 66 we see that $\beta_{1}$ increases at the beginning which facilitates taking larger steps at the early stage of training, while towards the end it decreases to give more emphasis on individual gradients. This decrease at the end is more severe in smaller tasks such as training on Shakespeare. Second, we observe that the interpolation factor of AVGrad, $\rho$, follows a three-phase pattern. In early iterations it increases, than it decreases and stays 0 until the middle of training where it starts to increase again. This suggests that MADA puts more importance on individual normalization term in the beginning and towards the end of training, and it puts more emphasis on the less noisy averaging term during the intermediate stages of training.
@@
Another interesting behavior we observe is when we initialize $\beta_{3}$ (which governs the weight of Adan) from a higher value such as $0.9$ (as in Figure 6), we see $\beta_{1}$ reaches 0.99 and stays constant, in contrast with Figures 66. The original paper Xie et al. (2023) suggests $\beta_{1}$ to be set a higher value such as $0.98$ (Note that $\beta_{1},\beta_{2},\beta_{3}$ in this work corresponds to $1-\beta_{1},1-\beta_{3},1-\beta_{2}$ in Xie et al. (2023)); remarkably, MADA initialized with $\beta_{3}=0.9$ automatically trains $\beta_{1}$ to be 0.99. In other words, MADA, when initialized from combination of Adan and AVGrad, automatically finds the $\beta_{1}$ that is suggested by the authors of (Xie et al., 2023).
@@
Experiments on Shakespeare dataset. On Shakespeare dataset we compare MADA to Adam and HyperAdam. We show that it results in significant increase in the performance across many initial $\beta_{1},\beta_{2}$ values. In Figure 7, we see that MADA results in better performance for the vast majority of initial $\beta_{1}$ and $\beta_{2}$ parameters, illustrating the robustness of our method. Note that in this figure we do not tune the hyper-hyper-parameters for individual initial $\beta$s.
Vision tasks. We also validated the performance of MADA on two computer vision models: 5-layer CNN and ResNet-9 on CIFAR-10 dataset and observe that MADA shows consistent improvement over our baselines (see test accuracy results below in Table 4).
Model MADA Adam HyperAdam SGD with Momentum
5-layer CNN 66.12 $\pm$ 0.14 $65.84\pm 0.11$ $65.80\pm 0.21$ $65.97\pm 0.94$
ResNet-9 93.79 $\pm$ 0.11 $93.73\pm 0.10$ $93.69\pm 0.06$ $92.60\pm 0.10$
@@ 6 Conclusion
In this paper, we propose an unified framework to parameterize the design space of a given collection of optimizers. Based on the framework, we develop MADA as a meta-adaptive optimizer that dynamically learns the coefficients of a parameterized optimizer during training. We further show, we can use MADA to dynamically learn the most suitable optimizer for a particular task during training using hyper-gradient descent. We employ MADA to train language and vision models and observe that it consistently outperforms Adam and other fixed optimizers both in terms of best validation performance, and in terms of robustness to initial hyper-parameters.
Limitations. The main limitation of our framework is the additional computational requirement and memory usage of the parameterized optimizer due to the maintained optimizer states. The additional computation can be broken down into two components: additional per-parameter computational steps during optimizer update, and the computation of hyper-gradients. Both can be shown to be negligible with respect to the computational requirements of overall training (see Appendix D). The additional memory usage arises from the fact that we need to maintain a larger optimizer state per parameter. To mitigate this, one can employ sharded data parallelism techniques (Rajbhandari et al., 2020), which shards the optimizer state across a large number of data-parallel devices. Another approach is to analyze the contributions of the constituent optimizers to the learning performance, and prune the ones that have little contribution.
@@ C.2 Training setup for the experiments
On OpenWebText, we use a global batch size of 480 sequences, cosine learning rate schedule with the peak learning rate of $6\times 10^{-4}$ ($1.5\times 10^{-4}$ for Lion) and the final learning rate of $1.5\times 10^{-5}$ (as chosen for Sophia algorithm in [Liu et al., 2023]). We use a weight decay of $0.1$ for all methods. We run the experiment for 100,000 iterations (first 2000 are warmup iterations) which corresponds to training on $\sim 492$B tokens. We use a weight decay parameter of $0.1$. On Shakespeare, we use a batch size of 64, cosine learning rate schedule with the peak learning rate of $10^{-3}$ and the final learning rate of $10^{-4}$. We run the experiment for 5,000 iterations (first 100 are warmup iterations). We run our experiments on AWS p5.48xlarge instances equipped with 8 NVIDIA H100 GPUs. To be able utilize multiple GPU’s we allreduced hyper-gradients across GPU’s.
@@
We initialize MADA with $\beta_{3}=0.9,\rho=0$ on top of Adam’s established $\beta_{1},\beta_{2}$ parameters for OpenWebText experiments as it resulted in a better validation loss. For Shakespeare experiments, we compare MADA against Adam over a grid of $(\beta_{1},\beta_{2})$ values, where in the case of MADA $(\beta_{1},\beta_{2})$ represents the initial values. In all experiments we initialize $\gamma=1,c=1$. For the OpenWebText experiments, we do not update hyper-parameters for the first 50 iterations for the sake of stability. For both experiments we use SGD as the hyper-parameter optimizer. On Shakespeare, we use $32.5e{-}$ 3 learning rate and $0.5$ momentum (we do not use momentum for $\gamma$) for learning the hyper-parameters. On OpenWebText we use a learning rate of $45e{-}$ 4 for training $\beta_{1},\beta_{2}$ and $11e{-}$ 1 for other hyper-parameters. Note that during the updates, we project every variable onto $[0,1]$ except for $\beta_{2}$ we project onto $[0.5,0.99]$ and for $\beta_{1},\beta_{3}$ we project onto $[0,0.99]$.
@@
Vision Tasks. For 5-layer model experiments, we use a CNN whose first two layers are convolutional layers with 6 and 16 output channels and 5 kernel size; and last 3 layers are fully connected layers with 120, 84, and 10 output dimensionality. We use ReLU activation in all layers except the last one and maxpool on the outputs convolutional layers. We use a constant learning rate of $31e{-}$ 3 for MADA, Adam, HyperAdam; and $21e{-}$ 2 learning rate and 0.9 momentum coefficient for SGD with momentum. We initialize MADA from Adam state, we set hyper lr for $\beta_{1},\beta_{2}$ to $41e{-}$ 4 and for the other variables to $21e{-}$ 2. We use batch size of 256 and train for 50 epochs. For ResNet-9 experiments, we use the model implementation from https://github.com/Moddy2024/ResNet-9. We use one cycle lr scheduler with $1e{-2}$ peak lr for MADA, Adam, HyperAdam and $1e{-1}$ peak lr for SGD with momentum. Again we initialize MADA from Adam, we set hyper lr for $\beta_{1}$ to $41e{-}$ 4, for $\beta_{2}$ to $41e{-}$ 4 and for the other variables to $21e{-}$ 2.. We use batch size of 400 and train for 50 epochs.

### v3

@@ 5.3 Results
Evolving hyper-parameters. We monitor the evolution of the interpolation coefficients during training of GPT-2 (125M) model on OpenWebText for various choices of initialization of $\beta_{1},\beta_{2}$ (Figures  7, 7, and 7).
@@
First, we observe that sub-optimal initialization of hyper-parameters (as in Figure 7) results in a more significant update of the coefficients. This suggests that default optimizer state with $\beta_{1,0}=0.9,\beta_{2,0}=0.95$ is close to a local minimum. A closer inspection yields insight into how MADA accelerates training. First, in Figures 77 we see that $\beta_{1}$ increases at the beginning which facilitates taking larger steps, while towards the end it decreases to give more emphasis on individual gradients. Second, we observe that the interpolation factor of AVGrad, $\rho$, follows a three-phase pattern. This suggests that MADA puts more importance on individual normalization term in the beginning and towards the end, and it puts more emphasis on the less noisy averaging term during the intermediate stages.
@@
Another interesting behavior we observe is when we initialize $\beta_{3}$ (which governs the weight of Adan) from a higher value such as $0.9$ (Figure 7), we see $\beta_{1}$ reaches 0.99 and stays constant, in contrast with Figures 77. Remarkably, MADA initialized with $\beta_{3}=0.9$ automatically trains $\beta_{1}$ to be 0.99; i.e. MADA, when initialized from combination of Adan and AVGrad, automatically finds the $\beta_{1}$ that is suggested by the authors of (Xie et al., 2023). Note that $\beta_{1},\beta_{2},\beta_{3}$ in this work correspond to $1-\beta_{1},1-\beta_{3},1-\beta_{2}$ in Xie et al. (2023).
@@
Shakespeare dataset. On Shakespeare dataset we compare MADA to Adam and HyperAdam. We show that it results in significant increase in the performance across many initial $\beta_{1},\beta_{2}$ values when training 11M parameter model. Particularly, in Figure 8, we see that MADA results in better performance for the vast majority of initial $\beta_{1}$ and $\beta_{2}$ parameters, illustrating robustness.
Fine-tuning GPT-2 XL on Shakespeare dataset. We also fine-tune the pre-trained GPT-2 XL (1.5B parameters) model on Shakespeare, for approximately 2 epochs, using MADA, Adam, and HyperAdam methods. Table 4 shows the resulting training loss values, where MADA results in a large improvement. Notably, HyperAdam achieves no gain over Adam.
Table 4: Training losses after 2 epochs of fine-tuning on Shakespeare dataset.
Method Training loss
MADA 0.255
Adam 0.276
HyperAdam 0.278
@@
Method 5-layer CNN ResNet-9
MADA 66.12 $\pm$ 0.14 93.79 $\pm$ 0.11
Adam $65.84\pm 0.11$ $93.73\pm 0.10$
HyperAdam $65.80\pm 0.21$ $93.69\pm 0.06$
Momentum SGD $65.97\pm 0.94$ $92.60\pm 0.10$
Vision tasks. We also validated the performance of MADA on two computer vision models: 5-layer CNN and ResNet-9 on CIFAR-10 dataset. We observe that MADA shows consistent improvement over our baselines (see test accuracy results below in Table 5).
@@ 6 Conclusion
In this paper we propose an approach to unify a set of optimizers into a parameterized formulation. We further show we can use MADA to dynamically learn the most suitable optimizer for a particular task during training using hyper-gradient descent. We employ MADA to train language models and observe that it consistently outperforms Adam and other fixed optimizers both in terms of best validation performance, and in terms of robustness to initial hyper-parameters.
Limitations. The main limitation of our framework is the additional computational requirement and memory usage of the parameterized optimizer due to the maintained optimizer states. The additional computation can be broken down into two components (per-parameter computational steps and computation of hyper-gradients); both can be shown to be negligible with respect to the computational requirements of overall training (see Appendix D, where we also provide the resulting memory use and iteration speed numbers). The additional memory usage arises from the fact that we need to maintain a larger optimizer state per parameter. To mitigate, one can employ sharded data parallelism techniques (Rajbhandari et al., 2020), which shards the optimizer state across a large number of data-parallel devices. Another approach is to analyze the contributions of the constituent optimizers to the learning performance, and prune the ones that have little contribution.
@@ C.2 Training setup for the experiments
On OpenWebText, we use a global batch size of 480 sequences, cosine learning rate schedule with the peak learning rate of $6\times 10^{-4}$ ($1.5\times 10^{-4}$ for Lion) and the final learning rate of $1.5\times 10^{-5}$ (as chosen for Sophia algorithm in [Liu et al., 2023]). For Lion we use the tuned learning rate from (Liu, 2023) for the same setting. For Adam we found that employing Sophia’s learning rate schedule results in better loss compared to employing Adam’s learning rate schedule in [Liu et al., 2023]. For our methods: AVGrad, MADA, MADA-FS, we directly employ Adam’s learning rate and learning schedule without any tuning since our goal is to demonstrate that MADA can be plugged in place of Adam without any changes in learning rate schedule.
@@
We run the experiment for 100,000 iterations (first 2000 are warmup iterations) which corresponds to training on $\sim 492$B tokens. We use a weight decay parameter of $0.1$. On Shakespeare, we use a batch size of 64, cosine learning rate schedule with the peak learning rate of $10^{-3}$ and the final learning rate of $10^{-4}$. We run the experiment for 5,000 iterations (first 100 are warmup iterations). We run our experiments on AWS p5.48xlarge instances equipped with 8 NVIDIA H100 GPUs. To be able utilize multiple GPU’s we allreduced hyper-gradients across GPU’s.
@@
We initialize MADA with $\beta_{3}=0.9,\rho=0$ on top of Adam’s established $\beta_{1},\beta_{2}$ parameters for OpenWebText experiments as it resulted in a better validation loss. For Shakespeare experiments, we compare MADA against Adam over a grid of $(\beta_{1},\beta_{2})$ values, where in the case of MADA $(\beta_{1},\beta_{2})$ represents the initial values. For the OpenWebText experiments, we do not update hyper-parameters for the first 50 iterations for the sake of stability. For both experiments we use SGD as the hyper-parameter optimizer. On Shakespeare, for 10M model we use $32.5e{-}$ 3 learning rate and $0.5$ momentum (we do not use momentum for $\gamma$) for learning the hyper-parameters. On OpenWebText (and for 1.5B model on Shakespeare) we use a learning rate of $45e{-}$ 4 for training $\beta_{1},\beta_{2}$ and $11e{-}$ 1 for other hyper-parameters.
@@
Vision Tasks. For 5-layer model experiments, we use a CNN whose first two layers are convolutional layers with 6 and 16 output channels and 5 kernel size; and last 3 layers are fully connected layers with 120, 84, and 10 output dimensionality. We use ReLU activation in all layers except the last one and maxpool on the outputs convolutional layers. We use a constant learning rate of $31e{-}$ 3 for MADA, Adam, HyperAdam; and $21e{-}$ 2 learning rate and 0.9 momentum coefficient for SGD with momentum. We initialize MADA from Adam state, we set hyper learning rate for $\beta_{1},\beta_{2}$ to $41e{-}$ 4 and for the other variables to $21e{-}$ 2. We use batch size of 256 and train for 50 epochs. For ResNet-9 experiments, we use the model implementation from https://github.com/Moddy2024/ResNet-9. We use one cycle learning rate scheduler with $1e{-2}$ peak learning rate for MADA, Adam, HyperAdam and $1e{-1}$ peak learning rate for SGD with momentum. Again we initialize MADA from Adam, we set hyper learning rate for $\beta_{1}$ to $41e{-}$ 4, for $\beta_{2}$ to $41e{-}$ 4 and for the other variables to $21e{-}$ 2. We use batch size of 400 and train for 50 epochs.